OpenStudy (anonymous):

y=sqrt(1+4sinx) (0,1) find the tangent to the curve at given point

OpenStudy (anonymous):

\[\huge y=\sqrt{1+4sinx}=(1+4sinx)^{\frac{1}{2}} \] \[\huge y'=\frac{1}{2}(1+4sinx)^{-\frac{1}{2}}\cdot (4cosx) \]

OpenStudy (anonymous):

now evaluate the derivative at x=0 to find the slope of the tangent line.

OpenStudy (anonymous):

so do you plug it in?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

\[\huge y'=\frac{1}{2}(1+4sinx)^{-\frac{1}{2}}\cdot (4cosx) \] \[\huge y'|_{x=0}=\frac{1}{2}(1+4sin(0))^{-\frac{1}{2}}\cdot (4cos(0)) \] \[\huge y'|_{x=0}=\frac{1}{2}(1+(0))^{-\frac{1}{2}}\cdot 4(1) \] \[\huge y'|_{x=0}=\frac{1}{2}\cdot 4 = 2\]

OpenStudy (anonymous):

k thats what i got

OpenStudy (anonymous):

y=2x+9

OpenStudy (anonymous):

now that you have a slope, m=2, and a point, (0, 1), put this into the point-slope form: \[\huge y-y_0=m(x-x_0) \]

OpenStudy (anonymous):

\[\huge y-y_0=m(x-x_0) \] \[\huge y-1=2(x-0) \] \[\huge y-1=2x \] \[\huge y=2x+1 \]

OpenStudy (anonymous):

ohhh ok!